OpenStudy (anonymous):

new in chemistry---In 3.45mg of boron there are how many of the following? grams of boron? moles of boron? boron atoms?

OpenStudy (anonymous):

grams of boron= are you sure its not 3.45 grams? because then it quite simply that, otherwise its 0.00345g of boron. moles of boron = mass/molar mass = 0.00345/10.81 = 0.000319148 moles boron atoms - 6.02x10^23x0.00345 = 2.076x10^21